# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ IV.3 Small-scale contributions
As for the ${1/N}$ case, let us first focus on the contributions from small scales. Indeed, Eq. (31) involves an infinite sum over the resonance numbers ${(k_{1},k_{2})}$, i.e. one ought to ensure that the limits ${k_{1},k_{2}\!\to\!+\infty}$ are meaningful.
@@
Figure 11: Same as Fig. 4, but for the ${1/N^{2}}$ Landau equation (Eq. 31) and profile $(2)$ (see Eq. 34). For small $k$, i.e. scales larger than $\epsilon$, we recover the expected scaling in $k^{-2}$.
From that figure, one can make a couple of observations: (i) For $k$ small enough, one finds ${\mathcal{R}_{2}(k)\!\propto\!k^{-2}}$. This is in agreement with the decay rate that can directly be inferred from Eq. (31), as detailed in Appendix F. (ii) As one increases the softening length, $\epsilon$, the contribution to the relaxation rate decays faster with $k$. As expected, introducing softening leads to the vanishing of the contribution from small scales. Overall, we find that Eq. (31) does not suffer from any small scale divergence, when applied to the dynamics of point vortices.
@@ IV.4 Relaxation rate
Building upon the previous case of profile (1), one could be worried that resonance broadening effects could be polluting the measurement from Fig. 12. In Fig. 13, we repeat the measurement of the relaxation rate, $\mathcal{R}_{2}$, for various values of $N$ and $q$.
Figure 13: Same as Fig. 13, but for profile (2). As one varies $N$ and/or $q$, the (rescaled) relaxation rate, $\mathcal{R}_{2}$ (Eq. 33), remains, in essence, unchanged.
@@ V Conclusions
In this paper, we investigated the long-term relaxation of long-range interacting point vortices in ${2D}$ hydrodynamics within axisymmetric configurations. We showed how the generic kinetic theory of inhomogeneous long-range interacting systems can readily be applied to point vortices. In that context, we emphasised the existence of two regimes of relaxation: (i) if the system’s mean frequency profile is non-monotonic, the distribution of point vortices relaxes through ${1/N}$ effects. In the limit of negligible collective amplification, this is governed by the ${1/N}$ inhomogeneous Landau equation (Eq. 10); (ii) if the mean frequency profile is monotonic, the system is then submitted to a kinetic blocking, which prevents ${1/N}$ effects from driving any relaxation. In that case, neglecting once again collective effects, the system’s relaxation is described by the ${1/N^{2}}$ inhomogeneous Landau equation (Eq. 31).
@@
${1/N^{2}}$ Balescu–Lenard equation. At order ${1/N}$, when collective effects cannot be neglected, the Landau equation becomes the Balescu–Lenard equation (Heyvaerts, 2010; Chavanis, 2012c). In the presence of a kinetic blocking, the ${1/N^{2}}$ version of the Balescu–Lenard equation is currently unknown. Although this is no easy challenge, it would be worthwhile to derive such a generalisation, to investigate the impact of collective effects in kinetic blockings.
@@ Appendix A Background profiles
with ${\Gamma_{\mathrm{tot}}^{\mathrm{b}}\!=\!\Gamma_{\mathrm{tot}}}$ the background total circulation. Here, $J_{\mathrm{b}}$ is the shift of the Lorentzian, while $\sigma_{\mathrm{b}}$ is its width. Using once again Eq. (36), one can compute explicitly the associated, unsoftened, potential. Although somewhat lengthy, this expression can be obtained using symbolic computing. In practice, in the $N$-body simulations, we imposed ${J_{\mathrm{b}}/J_{0}\!=\!\tfrac{1}{2}}$ and ${\sigma_{\mathrm{b}}/J_{0}\!=\!1}$, for which the generated frequency profile is non-monotonic in the vicinity of the active DF, see Fig. 2.
@@ Appendix C Linear response theory
[…]
@@
First, in Eq. (47), we note that both the resonance number, $k$, and the frequency, $\omega$, are external labels of the integral over ${\mathrm{d}J^{\prime\prime}}$. Let us therefore assume that we can decompose the action domain into a finite number of bins of length ${\Delta J}$. This provides us then with a list of action, ${\{J_{i}\}_{i}}$, around which each bin is centered. We now introduce the matrices, $\mathbf{U}_{k}$ and ${\mathbf{U}^{\mathrm{d}}_{k}(\omega)}$, via
@@
[…]
@@
Importantly, the present approach (i) does not require any potential/density basis elements; (ii) can readily be applied to softened pairwise interactions.
Nyquist contours. In Figs. 3 and 10, the Nyquist contours represent, in the complex plane, the values of the function ${\omega_{\mathrm{R}}\!\in\!\mathbb{R}\!\to\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)]\!=\!|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|}$. Phrased differently, these figures show the imaginary part of ${|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|}$ as a function of its real part, as one varies $\omega_{\mathrm{R}}$. In practice, (i) we used $500$ points to sample the action domain in Eq. (49), and (ii) we regularised Eq. (51) with the small positive imaginary part ${\eta\!=\!10^{-3}\,\Omega(J_{0})}$, so that ${1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)_{-}\!\to\!1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)}$. We checked that varying the number of sampling points or the value of $\eta$ had no impact on the plotted Nyquist curves.
Appendix D ${1/N}$ kinetic theory
${1/N}$ Landau equation. Following Chavanis (2012c), in the absence of collective effects, the inhomogeneous Landau equation (Eq. 10) satisfies
@@ Appendix D ${1/N}$ kinetic theory
Small-scale contributions. In Fig. 4, we present the respective contribution of small scale resonances, i.e. large $k$, to the ${1/N}$ Landau relaxation rate. Let us recover the main features of that figure directly from Eq. (10). For ${J^{\prime}\!\to\!J}$, Eq. (45b) gives ${U_{k}(J,J^{\prime})\!\simeq\!k^{-1}}$, so that Eq. (56) gives ${|\Lambda_{k}(\mathbf{J})|^{2}\simeq k^{-2}}$. When used in Eq. (15), we find from Eq. (14) that each scale contributes like $k^{-1}$. In practice, this scaling only holds for $k$ small enough, before the difference between the two resonating actions $J$ and $J_{1}$ leads a decay faster than $k^{-1}$ in Eq. (45b). Similarly, increasing the considered softening length accelerates, as expected, the decay of the small-scale contributions. These various trends are all recovered in Fig. 4.
@@
$\delta_{\mathrm{D}}\big{[}\Omega(J)\!-\!\Omega(J_{1})\big{]}=\sum_{J_{\star}}\frac{\delta_{\mathrm{D}}(J\!-\!J_{\star})}{|\mathrm{d}\Omega/\mathrm{d}J_{\star}|},$ (57)
where the $J_{\star}$ satisfy ${\Omega(J_{\star})\!=\!\Omega(J)}$. Used in Eq. (14), we get
$\frac{\partial F(J,t)}{\partial t}=\gamma\frac{\partial}{\partial J}\bigg{[}\sum_{J_{\star}}{}$ $\frac{\big{|}\Lambda_{\mathrm{tot}}(J,J_{\star})\big{|}^{2}}{|\mathrm{d}\Omega/\mathrm{d}J_{\star}|}$ (58)
[…]
We implemented this expression to obtain Fig. 5, limiting the sum over $k$ in Eq. (15) to ${|k|\!\leq\!100}$.
Computing the regularised prediction. To compute prediction from the “regularised” ${1/N}$ Landau equation (Eq. 25), we evaluated the integral over ${\mathrm{d}J_{1}}$ therein using a simple midpoint rule with 1 000 nodes within the interval ${J_{0}\!-\!\sigma_{0}\!\leq\!J_{1}\!\leq\!J_{0}\!+\!\sigma_{0}}$, limiting, once again, the sum over $k$ in Eq. (15) to ${|k|\!\leq\!100}$.
Inferring the regularisation time. In Fig. 8, we compare the value of nonlinear time, $T_{\mathrm{nl}}$ (Eq. 26), with the regularisation time, $T_{\mathrm{reg}}$, inferred from $N$-body simulations. To obtain this figure, we proceeded as follows. First, we computed the regularised relaxation rate from Eq. (25) for ${T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!10\!\times\!2^{n}}$ with ${0\!\leq\!n\!\leq\!7}$. Then, for each such prediction, we performed a linear fit of the relaxation rate within the domain ${1.022\!\leq\!J/J_{0}\!\leq\!1.042}$, the action domain centered around the change of sign of the relaxation rate. These measurements are presented in Fig. 14.
@@ Appendix F ${1/N^{2}}$ kinetic theory
with the coupling functions, $\mathcal{U}^{(1)}$ and $\mathcal{U}^{(2)}$, defined in Appendix B of Fouvry (2022). In Eq. (31), the sum over the resonance vectors ${(k_{1},k_{2})}$ is restricted to those such that $k_{1}$, $k_{2}$, as well ${k_{1}\!+\!k_{2}}$ are all non-zero. Importantly, we emphasise that ${|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}}$ is independent of $N$ and $q$.
Small-scale contributions. In Fig. 11, we present the contributions from small scales to the ${1/N^{2}}$ relaxation rate. Starting from Eq. (31), let us place ourselves in the limit where the three actions, $\mathbf{J}$, involved in the three-body resonance condition are close to one another, and assume that ${k_{1},k_{2}\!\simeq\!k}$. Then, from eq. (B1) in Fouvry (2022), we find that ${\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!k^{2}U_{k}^{2}}$, where we assumed that computing gradients of the form ${\partial_{J}U_{k_{1}}(J,J^{\prime})}$ does not lead to any additional scaling dependence with respect to $k$. Imposing ${U_{k}\!\simeq\!k^{-1}}$ following Eq. (45b), we obtain ${\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!1}$ as a function of $k$. Similarly, starting from eq. (B2) in Fouvry (2022), we find ${\mathcal{U}^{(2)}_{k_{1}k_{2}}\!\simeq\!k^{-1}}$. With these two scalings, we find from Eq. (62) that ${|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}\!\simeq\!k^{-4}}$. Injecting this dependence in Eq. (31), and using the definition from Eq. (35), we get ${\mathcal{R}_{2}(J,k_{1},k_{2})\!\simeq\!k^{-3}}$. Finally, we may perform the sum over ${k_{1},k_{2}}$ in Eq. (35), with the constraint ${\mathrm{Max}[|k_{1}|,|k_{2}|]\!=\!k}$, to obtain ${\mathcal{R}_{2}(J,k)\!\simeq\!k^{-2}}$. This is the scaling that is indeed recovered in Fig. 11. As expected, in that figure, we also recover that softening accelerates the decay of the contributions stemming from small scales.
@@ Appendix G Numerical simulations
Time integration. Given that the Hamiltonian from Eq. (8) is not under a straightforward separable form, we perform the time integration using a standard explicit Runge–Kutta method. Yet, since both the external potential and the pairwise interaction potentials are smooth, we may resort to high-order methods. This allows for the use of a large individual timestep making long-time integrations easier to perform. In that view, we used the 15-stage 9th-order “more efficient” Vern9 method available through the library OrdinaryDiffEq.jl.
@@
[…]
For profile (2), we fixed the timestep to ${h/T_{\mathrm{dyn}}\!\simeq\!1.90\!\times\!10^{-2}}$. Each simulation was integrated up to ${T_{\mathrm{max}}/T_{\mathrm{dyn}}\!\simeq\!3.040\!\times\!10^{6}\,(0.0005/q)^{4}(N/200)^{2}}$, to match the scaling from Eq. (32). Similarly, we made 2 000 dumps of each realisation as the integration was performed. For that profile, the final relative errors on $E_{\mathrm{tot}}$ and $L_{\mathrm{tot}}$ were of order $10^{-9}$. In practice, the case ${(N,q)\!=\!(200,0.0005)}$ presented in Fig. 12 required ${\!\sim\!21\mathrm{h}}$ of computation time, per realisation, on a single core.
[…]
Figure 15: Time evolution of the number of particles, ${N(\!<\!J,t)}$, in numerical simulations, when averaged over all available realisations, for profile (1) [top] and profile (2) [bottom]. Different colors correspond to different actions $J$. The linear dependence of these time series is the sign of the onset of relaxation.
In that figure, we note that the relative change in the fraction of particles is of order $10^{-3}$, i.e. we purposely limited our measurements to the onset of relaxation. Once the average time series ${t\!\mapsto\!\langle N(\!<\!J,t)\rangle}$ determined, we performed a linear fit of the form
@@
Finally, we use the slope $\beta$ in Eqs. (17) and (33) to estimate the relaxation rates. Given the computing cost associated with each such realisation, the number of independent realisations, as well as the large number of values of ${(N,q)}$ considered, obtaining Fig. 8 was no light numerical endeavour. Finally, to obtain Fig. 9, we binned the action domain in bins of width ${10^{-2}\,J_{0}}$.

### v2

@@ IV.3 Small-scale contributions
As for the 1/N{1/N} case, let us first focus on the contributions from small scales. Indeed, Eq. (40) involves an infinite sum over the resonance numbers $(k_{1},k_{2})$}, i.e. one ought to ensure that the limits $k_{1},k_{2}\!\to\!+\infty$} are meaningful.
@@
Figure 12: Same as Fig. 4, but for the $1/N^{2}$} Landau equation (Eq. 40) and profile $(2)$ (Eq. 43). For small $k$, i.e. scales larger than $\epsilon$, we recover the expected scaling in $k^{-2}$.
From that figure, one can make a couple of observations: (i) For $k$ small enough, one finds $\mathcal{R}_{2}(J,k)\!\propto\!k^{-2}$}. This is in agreement with the decay rate that can directly be inferred from Eq. (40), as detailed in Appendix F. (ii) As one increases the softening length, $\epsilon$, the contribution to the relaxation rate decays faster with $k$. As expected, introducing softening leads to the vanishing of the contribution from small scales. Overall, we find that Eq. (40) does not suffer from any small scale divergence (even in the absence of softening), when applied to the dynamics of point vortices.
@@ IV.4 Relaxation rate
Building upon the previous case of profile (1), one could be worried that resonance broadening effects could be polluting the measurement from Fig. 13. In Fig. 14, we repeat the measurement of the relaxation rate, $\mathcal{R}_{2}(J)$}, for various values of $N$ and $q$.
Figure 14: Same as Fig. 14, but for profile (2). As one varies NN and/or $q$, the (rescaled) relaxation rate, $\mathcal{R}_{2}(J)$} (Eq. 42), remains, in essence, unchanged.
@@ V Conclusions
In this paper, we investigated the long-term relaxation of long-range interacting point vortices in 2D{2D} hydrodynamics within axisymmetric configurations. We showed how the generic kinetic theory of inhomogeneous long-range interacting systems can readily be applied to point vortices. In that context, we emphasised the existence of two regimes of relaxation: (i) if the system’s mean frequency profile is non-monotonic, the distribution of point vortices relaxes through 1/N{1/N} effects. In the limit of negligible collective amplification, this is governed by the 1/N{1/N} inhomogeneous Landau equation (Eq. 10); (ii) if the mean frequency profile is monotonic, the system is then submitted to a kinetic blocking, which prevents 1/N{1/N} effects from driving any relaxation. In that case, neglecting once again collective effects, the system’s relaxation is described by the $1/N^{2}$} inhomogeneous Landau equation (Eq. 40).
@@
$1/N^{2}$} Balescu–Lenard equation. At order 1/N{1/N}, when collective effects cannot be neglected, the Landau equation becomes the Balescu–Lenard equation (Heyvaerts, 2010; Chavanis, 2012c). In the presence of a kinetic blocking, the $1/N^{2}$} version of the Balescu–Lenard equation is currently unknown. Although this is no easy challenge, it would be worthwhile to derive such a generalisation, to investigate the impact of collective effects in kinetic blockings.
@@ Appendix A Background profiles
with $\Gamma_{\mathrm{tot}}^{\mathrm{b}}\!=\!\Gamma_{\mathrm{tot}}$} the background total circulation. Here, $J_{\mathrm{b}}$ is the shift of the background DF, while $\sigma_{\mathrm{b}}$ is its width. Using once again Eq. (45), one can compute explicitly the associated, unsoftened, potential. Although somewhat lengthy, this expression can be obtained using symbolic computing. In practice, in the $N$-body simulations, we imposed $J_{\mathrm{b}}/J_{0}\!=\!\tfrac{1}{2}$} and $\sigma_{\mathrm{b}}/J_{0}\!=\!1$}, for which the generated frequency profile is non-monotonic in the vicinity of the active DF, see Fig. 2.
@@ Appendix C Linear response theory
$+{}$  $\,2\pi\!\!\int\!\!\mathrm{d}J^{\prime\prime}\frac{k\partial F/\partial J^{\prime\prime}}{(k\Omega^{\prime\prime}-\omega)_{-}}\,U_{k}(J,J^{\prime\prime})\,U^{\mathrm{d}}_{k}(J^{\prime\prime},J;\omega),$
@@
First, in Eq. (56), we note that both the resonance number, $k$, and the frequency, $\omega$, are external labels of the integral over $\mathrm{d}J^{\prime\prime}$}. Let us therefore assume that we can decompose the action domain into a finite number of bins of length $\Delta J$}. This provides us then with a list of action, $\{J_{i}\}_{i}$}, around which each bin is centred. We now introduce the matrices, $\mathbf{U}_{k}$ and $\mathbf{U}^{\mathrm{d}}_{k}(\omega)$}, via
@@
$\big(\mathbf{M}_{k}(\omega)\big)_{ij}=\delta_{ij}\,2\pi\,\Delta J\,\frac{k\partial F/\partial J_{i}}{(k\Omega(J_{i})-\omega)_{-}}.$ (60)
@@
Importantly, the present approach (i) does not require any potential/density basis elements; (ii) can readily be applied to softened pairwise interactions. Since the approach from this Appendix directly computes the dressed coupling coefficients, $U^{\mathrm{d}}$, it can readily be used to predict the BL relaxation rate (Heyvaerts, 2010; Chavanis, 2012c). This provides an alternative to the traditional basis method (Kalnajs, 1976) used, for example, in the BL predictions of Roule et al. (2022).
Nyquist contours. In Figs. 3 and 11, the Nyquist contours represent, in the complex plane, the values of the function $\omega_{\mathrm{R}}\!\in\!\mathbb{R}\!\to\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)]\!=\!|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|$}. Phrased differently, these figures show the imaginary part of $|\mathbf{E}_{k}(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)|$} as a function of its real part, as one varies $\omega_{\mathrm{R}}$. In practice, (i) we used $500$ points to sample the action domain in Eq. (58), and (ii) we regularised Eq. (60) with the small positive imaginary part $\eta\!=\!10^{-3}\,\Omega(J_{0})$}, so that $1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)_{-}\!\to\!1/(\omega_{\mathrm{R}}\!+\!\mathrm{i}\eta)$}. We checked that varying the number of sampling points or the value of $\eta$ had no impact on the plotted Nyquist curves.
Appendix D 1/N{1/N} kinetic theory
1/N{1/N} Landau equation. Following Chavanis (2012c), in the absence of collective effects, the inhomogeneous Landau equation (Eq. 10) satisfies
@@
Small-scale contributions. In Fig. 4, we present the respective contribution of small scale resonances, i.e. large $k$, to the 1/N{1/N} Landau relaxation rate. Let us recover the main features of that figure directly from Eq. (10). For $J^{\prime}\!\to\!J$}, Eq. (52b) gives $U_{k}(J,J^{\prime})\!\simeq\!k^{-1}$}, so that Eq. (65) gives $|\Lambda_{k}(\mathbf{J})|^{2}\simeq k^{-2}$}. When used in Eq. (15), we find from Eq. (14) that each scale contributes like $k^{-1}$. In practice, this scaling of the relaxation rate only holds for $k$ small enough, before the difference between the two resonating actions $J$ and $J_{1}$ leads a decay faster than $k^{-1}$ in Eq. (52b). For the diffusion coefficient from Eq. (17), because local resonances always contribute to it, i.e. $J_{1}\!=\!J$} in Eq. (17), the exact match of the two resonating actions allows for the scaling in $k^{-1}$ to hold throughout, as visible in Eq. (52b). This is responsible for a small-scale logarithmic divergence of the diffusion coefficient, in systems driven by the unsoftened interaction from Eq. (3) (see Appendix H). As could have been expected, increasing the considered softening length accelerates the decay of the small-scale contributions. These various trends are further discussed in Section III.3 and recovered in Fig. 4.
@@
$\delta_{\mathrm{D}}\big[\Omega(J)\!-\!\Omega(J_{1})\big]=\sum_{J_{\mathrm{r}}}\frac{\delta_{\mathrm{D}}(J\!-\!J_{\mathrm{r}})}{|\mathrm{d}\Omega/\mathrm{d}J_{\mathrm{r}}|},$ (66)
where the $J_{\mathrm{r}}$ satisfy $\Omega(J_{\mathrm{r}})\!=\!\Omega(J)$}. Used in Eq. (14), we get
$\frac{\partial F(J,t)}{\partial t}=\gamma\frac{\partial}{\partial J}\bigg[\sum_{J_{\mathrm{r}}}{}$ $\frac{\big|\Lambda_{\mathrm{tot}}(J,J_{\mathrm{r}})\big|^{2}}{|\mathrm{d}\Omega/\mathrm{d}J_{\mathrm{r}}|}$ (67)
$\times{}$ $\bigg(F(J_{\mathrm{r}})\frac{\partial F}{\partial J}\!-\!F(J)\frac{\partial F}{\partial J_{\mathrm{r}}}\bigg)\bigg].$
We implemented this expression to obtain Fig. 5, limiting the sum over $k$ in Eq. (15) to $|k|\!\leq\!100$}. We used the same approach to compute the 1/N{1/N} Landau diffusion coefficient from Eq. (17).
Computing the regularised prediction. To compute prediction from the “regularised” 1/N{1/N} Landau equation (Eq. 32), we evaluated the integral over $\mathrm{d}J_{1}$} therein using a simple midpoint rule with 1 000 nodes within the interval $J_{0}\!-\!\sigma_{0}\!\leq\!J_{1}\!\leq\!J_{0}\!+\!\sigma_{0}$}, limiting, once again, the sum over $k$ in Eq. (15) to $|k|\!\leq\!100$}. We proceeded similarly to evaluate the regularised diffusion coefficient from Eq. (33).
Inferring the regularisation time. In Fig. 9, we compare the value of the nonlinear time, $T_{\mathrm{nl}}$ (Eq. 35), with the regularisation time, $T_{\mathrm{reg}}$, inferred from $N$-body simulations. To obtain this figure, we proceeded as follows. First, we computed the regularised relaxation rate from Eq. (32) for $T_{\mathrm{reg}}/T_{\mathrm{dyn}}\!=\!10\!\times\!2^{n}$} with $0\!\leq\!n\!\leq\!7$}. Then, for each such prediction, we performed a linear fit of the relaxation rate within the domain $1.022\!\leq\!J/J_{0}\!\leq\!1.042$}, the action domain centred around the change of sign of the relaxation rate. These measurements are presented in Fig. 15.
@@ Appendix F $1/N^{2}$} kinetic theory
with the coupling functions, $\mathcal{U}^{(1)}$ and $\mathcal{U}^{(2)}$, defined in Appendix B of Fouvry (2022). In Eq. (40), the sum over the resonance vectors $(k_{1},k_{2})$} is restricted to those such that $k_{1}$, $k_{2}$, as well as $k_{1}\!+\!k_{2}$} are all non-zero. Importantly, we emphasise that $|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}$} is independent of $N$ and $q$.
Small-scale contributions. In Fig. 12, we present the contributions from small scales to the $1/N^{2}$} relaxation rate. Starting from Eq. (40), let us place ourselves in the limit where the three actions, $\mathbf{J}$, involved in the three-body resonance condition are close to one another, and assume that $k_{1},k_{2}\!\simeq\!k$}. Then, from eq. (B1) in Fouvry (2022), we find that $\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!k^{2}U_{k}^{2}$}, where we assumed that computing gradients of the form $\partial_{J}U_{k_{1}}(J,J^{\prime})$} does not lead to any additional scaling dependence with respect to $k$. Assuming local resonances, we may follow Eq. (52b) and impose the scaling $U_{k}\!\simeq\!k^{-1}$}. We then obtain $\mathcal{U}^{(1)}_{k_{1}k_{2}}\!\simeq\!1$} as a function of $k$. Similarly, starting from eq. (B2) in Fouvry (2022), we find $\mathcal{U}^{(2)}_{k_{1}k_{2}}\!\simeq\!k^{-1}$}. With these two scalings, we find from Eq. (71) that $|\Lambda_{k_{1}k_{2}}(\mathbf{J})|^{2}\!\simeq\!k^{-4}$}. Injecting this dependence in Eq. (40), and using the definition from Eq. (44), we get $\mathcal{R}_{2}(J,k_{1},k_{2})\!\simeq\!k^{-3}$}. Finally, we may perform the sum over $k_{1},k_{2}$} in Eq. (44), with the constraint $\mathrm{Max}[|k_{1}|,|k_{2}|]\!=\!k$}, to obtain $\mathcal{R}_{2}(J,k)\!\simeq\!k^{-2}$}. This is the scaling that is indeed recovered in Fig. 12. As expected, in that figure, we also recover that softening accelerates the decay of the contributions stemming from small scales.
@@ Appendix G Numerical simulations
Time integration. The Hamiltonian from Eq. (8) is not under a separable form. As such, one cannot readily use standard splitting methods to devise symplectic integration (see, e.g. Hairer et al., 2006). This could be circumvented by appropriately “duplicating phase space” (see, e.g., Tao, 2016). Though, such symplectic approaches typically require one to carefully tune some “binding constant” between the two copies of the system’s Hamiltonian. For a given problem, the optimal value of this ad hoc constant, noted $\omega$ in Tao (2016), can depend, among other things, on the value of $N$ or the considered active fraction, $q$. In the present context, since we are exploring a large range of values in $N$, and since both the external potential and the pairwise interaction potentials are smooth, it appears more convenient to resort to more traditional high-order explicit Runge–Kutta methods. In particular, this allows for the use of a large individual timestep making long-time integrations easier to perform. This also removes the need for any additional control parameter to be calibrated. In that view, we used the 15-stage 9th-order “more efficient” Vern9 method available through the library OrdinaryDiffEq.jl.
@@
For profile (1), we fixed the timestep to $h/T_{\mathrm{dyn}}\!\simeq\!1.41\!\times\!10^{-2}$}. Each simulation was integrated up to $T_{\mathrm{max}}/T_{\mathrm{dyn}}\!\simeq\!1.013\!\times\!10^{3}\,(0.0001/q)^{2}\,(N/2000)$}, to match the scaling of the relaxation time expected from Eq. (19). We made 2 000 dumps of each realisation as the integration was performed. The final relative errors on $E_{\mathrm{tot}}$ and $L_{\mathrm{tot}}$ were typically of order $10^{-13}$, as illustrated in Fig. 16.
Figure 16: Time evolution of the relative error in $E_{\mathrm{tot}}$ and $L_{\mathrm{tot}}$ for profile (1) (top panel) and profile (2) (bottom panel) during the $N$-body integration. The fillings illustrate the spread between the 16% and 84% level lines over the available realisations.
In that same figure, we recover that our explicit integration scheme leads to a linear growth in time of the error in the global $N$-body invariants. In practice, the case (N,q)=(2000,0.0001){(N,q)\!=\!(2000,0.0001)} presented in Fig. 5 required $\sim\!40\mathrm{min}$} of computation time per realisation, on a single core. Of course, simulating the case (N,q)=(8 000,0.0001){(N,q)\!=\!(8\,000,0.0001)}, explored in Fig. 9, was even more numerically costly, and required $\sim\!36\mathrm{h}$} of computation time per realisation.
For profile (2), we fixed the timestep to $h/T_{\mathrm{dyn}}\!\simeq\!1.90\!\times\!10^{-2}$}. Each simulation was integrated up to $T_{\mathrm{max}}/T_{\mathrm{dyn}}\!\simeq\!3.040\!\times\!10^{6}\,(0.0005/q)^{4}(N/200)^{2}$}, to match the scaling from Eq. (41). Similarly, we made 2 000 dumps of each realisation as the integration was performed. For that profile, the final relative errors on $E_{\mathrm{tot}}$ and $L_{\mathrm{tot}}$ were of order $10^{-9}$, as illustrated in Fig. 16. In practice, the case (N,q)=(200,0.0005){(N,q)\!=\!(200,0.0005)} presented in Fig. 13 required $\sim\!21\mathrm{h}$} of computation time, per realisation, on a single core.
Measuring relaxation. To obtain the numerical results presented throughout the main text, we proceeded as follows. For each realisation, we sliced the action domain, $0.8\!\leq\!J/J_{0}\!\leq\!1.2$} in 50 locations, uniformly separated. Then, for each such action, and each dump, we computed the instantaneous number of particles with action smaller than the considered action. Doing so, for each realisation and each considered action, we had at our disposal a time series of $t\!\mapsto\!N(\!<\!J,t)$}, as introduced in Eq. (20a). For each value of (N,q){(N,q)} considered, we ran a total of $N_{\mathrm{real}}\!=\!4\,096$} independent realisations in order to estimate ensemble averages. In Fig. 17, we illustrate the typical change in the number of particles, $\Delta N(\!<\!J,t)\!=\!N(\!<\!J,t)\!-\!N(\!<\!J,t\!=\!0)$}, when averaged over all available realisations.
Figure 17: Time evolution of the number of particles, N(<J,t){N(\!<\!J,t)}, in numerical simulations, when averaged over all available realisations, for profile (1) [top] and profile (2) [bottom]. Different colours correspond to different actions $J$. The linear dependence of these time series is the sign of the onset of relaxation.
In that figure, we note that the relative change in the fraction of particles is of order $10^{-3}$, i.e. we purposely limited our measurements to the onset of relaxation. Nonetheless, such a relative change is much larger than the typical relative error in the $N$-body invariants (see Fig. 16). This strengthens our confidence in the appropriate convergence of the $N$-body measurements of the relaxation rate.
Once the average time series $t\!\mapsto\!\langle N(\!<\!J,t)\rangle$} determined, we performed a linear fit of the form
@@
Finally, we use the slope $\beta$ in Eqs. (20a) and (42) to estimate the relaxation rates. Given the computing cost associated with each such realisation, the number of independent realisations, as well as the large number of values of (N,q){(N,q)} considered, obtaining Fig. 9 was no light numerical endeavour. Finally, to obtain Fig. 10, we binned the action domain in bins of width $10^{-2}\,J_{0}$}.